### v4

@@ 3.2.4 Each method has yet to reach optimum accuracy
Mejri and Ben-Othman (2014), The proposed fix is a detection system that uses a statistical method, linear regression, and a watchdog to discover greedy behaviour in the MAC layer passively. A watchdog monitors the degree to which active nodes’ access periods correlate as part of the suggested approach. The network is under assault if the correlation coefficient is close to 1 but not precisely 1, and the linear regression straight-line slope is also not close to 1. This can happen when the correlation coefficient is near one but not quite one or when the slope is very close to 1 but not quite 1. The developed program does this by monitoring three metrics: the interval between transmissions, the total time spent transferring data, and the number of connect attempts made by a node. Then, with the aid of the trace file created by SUMO, the solutions’ functioning is analysed in real-time. -3. The SUMO scenario is based on a city map with labels and signals. Researchers first simulate the network using nodes that behave correctly to ensure the linear regression method is reliable. After that, four greedy automobiles are produced by injecting greedy nodes one at a time. To detect greedy behaviour, this approach takes 1.3 seconds for 1 node, 1.9 seconds for 2 nodes, 3.1 seconds for 3 nodes, and 7.9 seconds for 4 nodes. Ali Alheeti et al. (2015) The solution proposed in the paper is an ANN-based ill-use and anomaly intrusion detection system that can identify DoS assaults. The authors designed the solution by first developing a mobility scenario within SUMO. For ns-2 to read the files SUMO created, a conversion was performed using Mobility Vehicle (MOVE). In the end, ns-2 was utilised to imitate the vehicles’ communication ability. When considering the mobility and traffic situation, the authors used the urban mobility of Manhattan as their inspiration. The simulation takes two minutes and fifty seconds, and it involves cars and six Road roadside units (RSUs). A CBR (Constant Bit Rate) traffic that conveys User Datagram Protocol (UDP) packets is being performed by the vehicles. Only one of the automobiles intentionally tries to cause harm. It will discard the packets rather than send them to the destination. The solution built was accurate 85.02% of the time when classifying normal behaviours and 98.45% when classifying deviant behaviours.
@@
Gruebler et al. (2015) The authors advocate using a Proportional Overlapping Scores (POS) approach to reduce the number of features taken from the trace file. These are fed into an artificial neural network (ANN) to teach it to classify data. After that, the ns-2 trace files that were created were employed in the POS algorithms so that the appropriate characteristics could be extracted. Next, Fuzzification is applied to the dataset’s data to prevent classification difficulties. Finally, the dataset is classified with the help of a Feed-Forward Neural Network (FFNN) using the IDS created for this solution. Sixty thousand records from the dataset were used, with half designated for training, another quarter for testing, and the remaining quarter for validation (25%). This dataset includes both benign and harmful activity in equal measure. The malicious behaviour vehicle was designed as a Black Hole attack, in which the malicious vehicle would not forward any received packets but dump them instead. Ultimately, the intrusion detection system was evaluated using anomaly and misuse detection. Regarding identifying misuse, it acquired a classification of 99.89 % for normal conduct and 99.80% for deviant behaviour. Regarding detecting anomalies, the findings for typical and aberrant behaviours came in at 99.87 and 99.72%, respectively.
@@
Sedjelmaci and Senouci (2015) This paper presents a cluster-based intrusion detection system (IDS) to defend a network against selective forwarding, black holes, wormholes, packet repetition, resource depletion, and Sybil attacks. The technique that has been presented uses several different detection managers that operate on three different levels: cluster associate, cluster head, and RSU. The cars’ speeds determine how they should be organised into clusters. Cluster connection and the assurance of data safety are the criteria utilised in selecting the cluster heads. When social behaviour is introduced, the connectivity inside a cluster is improved. The IDS architecture comprises two primary detection schemes and a decision scheme. Local Intrusion Detection System and Global Intrusion Detection System are the detection systems run at cluster members and heads. The name of the decision-making scheme is Global Decision System, which is hosted on the RSUs. Because of this, the system can monitor several organisations and identify threats on multiple levels. In addition, each level can perform a unique set of algorithms and detecting approaches, which evaluate a unique set of characteristics. The Global Decision System will provide the collective standing of each car, which the cluster head will then send. The Worldwide Decision Scheme will then calculate the trust level of every vehicle. To evaluate the proposed system, ns-3 was used as the network simulator, while SUMO was used to mimic the vehicle’s movement. The IDS was evaluated concerning the detection rate, the false positive rate, and the detection time. The various assaults were put through their paces using a range of vehicle counts during testing. The percentage of hostile cars always equals 45% of the total. The provided findings range from 92% to 100%, although this depends on the total number of vehicles and the type of attack conducted.
@@
Ali Alheeti et al. (2016) proposed A FFNN and SVM-based intrusion detection system is suggested in this paper. A methodical reaction is also meant to protect automobiles if harmful conduct is identified. A dataset composed of SUMO and ns-2 is utilised to train the IDS, and the features extracted from the trace file are used as training data. The POS algorithm has reduced these characteristics from 21 to 15. The collection consists of 30,000 entries that differentiate between appropriate and inappropriate behaviours. Fuzzification is performed on this dataset before it is used for classification. The writers separate the dataset into three categories: validation, test, and training. Generating grey-hole attacks involves picking malicious vehicles that drop packets at arbitrary intervals throughout the attack. The Ad Hoc On-Demand Distance Vector (AODV) protocol needs to be modified so that rushing assault can be carried out successfully. In SVM, the accuracy values achieved through simulation were 99.93% for normal behaviour and 99.64% for aberrant conduct. When FFNN was used, the findings showed an accuracy of 99.82% for normal behaviour and 98.86% accuracy for deviant conduct, respectively.
@@ 25 Features:
Alheeti et al. (2017) Using Linear and Quadratic Discriminant Analysis, the authors present an intelligent IDS to guard against assaults, namely DoS and Black Hole attacks. To begin, evil actions are modelled and then emulated. Next, the authors make several adjustments to the AODV protocol to produce the DoS attack. In this scenario, denial-of-service attacks are carried out by dropping packets. After that, multiple mobilities and traffic setups were developed. SUMO and MOVE produced a realistic setting of both abnormal and typical behaviour. In addition to that, ns-2 was utilised to imitate various forms of communication. The dataset is derived from the ns-2 trace file, and all 21 characteristics are preserved throughout the process. The data is given a fuzzy appearance before being put through the test and the training. The findings of the Linear Discriminant Analysis indicate a detection rate of 86.44%, whereas the Quadratic Discriminant Analysis results indicate a detection rate of 85.67%.
@@
Sharma and Kaul (2018) The authors suggest a novel technique for selecting a steady Cluster Head, and it uses Hybrid Fuzzy Multi-standards. After that, an ML-based intrusion detection system that uses SVM is utilised to identify malicious activities. The utilisation of a Dolphin Swarm Algorithm results in an improvement to the SVM-based IDS detection capabilities. This program uses dolphin swarms’ hunting and preying behaviour to locate and isolate misbehaving nodes in a network. The future method is evaluated through the use of simulation. Matlab and NetSim simulate the network, while SUMO models the traffic. During the simulation, a range of node densities from 50 to 300 are examined, with a maximum of 45% of the cars acting as attackers. The suggested intrusion detection system has a detection rate greater than 98%, selective forwarding, and wormhole attack; however, this percentage can vary based on the number of vehicles.
@@
Kumar et al. (2015) offered a cloud intrusion detection system founded on clustering, and Learning Automate (LA) was presented. This method can potentially improve detection activity. The LA determines the leadership of clusters based on factors such as the total comparative velocity and the connectedness grade. The next stage, safeguarding the data, will begin after this procedure. In this process stage, an automaton will validate messages using the HMAC algorithm. When utilising the method that has been suggested, around 93% of harmful actions are recognised, and when mobility is added to leadership formation, the percentage of false positives is reduced. In addition to that, this technique is flexible enough to adapt to any alterations made to the network’s nodes.
@@
Shu et al. (2021) presented an IDS for VANETs by implementing a Software-Defined Networking (SDN)-Based controller on each base station to differentiate between regular network traffic and attack network flows. Using the total network flow data, they trained several SDN controllers for the full VANET using GAN. This IDS approach permits distributed SDN controllers to detect their sub-network flows directly, and it can reduce high processing overheads. The author tested their algorithm using criteria along with precision, Accuracy, F1-Score, recall, and area under the curve (AUC), where maximum accuracy is 98.4%.
@@
Zeng et al. (2019) Researchers have proposed a deep learning-based intrusion detection method incorporating CNN and LSTM. The proposed model’s efficiency was compared to previously developed ML methods using a public dataset and an NS3 simulated VANET dataset with 28236 entries. The recommended model earned an F1 score ranging from 0.91 to 0.99.
@@
Gad et al. (2021) discussed ML-based IDS techniques in VANET that are impressive. Researchers used SMOTE, chi2, and SMOTE+chi2 to eliminate the database imbalance problem in the feature engineering phase, and the XGBoost method’s overall score is better than others. In the Training-Testing phase, researchers used Logistic regression (LR), naive Bayes (NB), decision tree (DT), support vector machine (SVM), k-nearest neighbour (kNN), Random Forest (RF), AdaBoost, and XGBoost. This experiment could be better by using the Ensemble Sacking classification to increase overall performance, which this paper wants to address.
Türkoğlu et al. (2022) used the MRMR feature selection algorithm to get the most correlated features for attack detection. First, scholars split the dataset in 3 ways: Training, Validation, and Test Data. After that, researchers used Bayesian Hyperparameter Optimization with training and validation data. However, 3-way splitting (60%, 20% and 20%) was not mandatory for tuning. Instead, the same training data could be used for tuning purposes. That way, 70% or 80% of data could be used for training the database, and 30% or 20% of data will be used for testing purposes to increase the overall accuracy.
Aneja et al. (2018) done proper research using ANN (feed-forward network) with backpropagation technique. This backpropagation operates by scheming the gradient at every layer, known as the local gradient. Then, the weights and bias values are updated according to Levenberg-Marquardt optimisation. Moreover, a Genetic algorithm is used for feature selection.
Alsarhan et al. (2021) use an optimised Support Vector Machine (SVM)-based Intrusion Detection System (IDS) for Vehicular Ad Hoc Networks (VANETs), employing Genetic Algorithm (GA), Particle Swarm Optimisation (PSO) and Ant Colony Optimization (ACO) to enhance accuracy. GA outperformed other algorithms, demonstrating improved detection rates and minimised false positives. A comprehensive evaluation validates the approach on the NSL-KDD dataset, providing significant insights into VANET security.
@@ 4 Technical Implementation
The VeReMi dataset was created to test the effectiveness of VANET misbehaviour detection systems (vehicular networks). This database contains onboard unit message logs created from a simulation environment and tagged ground truth. The collection contains malicious messages that are meant to cause erroneous application behaviour, which misbehaviour detection techniques are designed to stop. The initial dataset includes five types of position falsification attacks. van der Heijden (2018). This data was generated by van der Heijden (2018), and It has been collected from Sharma (2021).
The CICIDS2017 intrusion detection dataset was used in this study because it comprises the most recent network attack scenarios, including those that are often employed in VANETs. Furthermore, no real security dataset with recent VANET threats is publicly accessible. In contrast, consider the fact that most passenger vehicles are linked to networks for entertainment as well as for transportation. To ensure a good user experience, IDS must identify malware activity from regular network traffic. The benign and common assaults in the CICIDS2017 dataset closely mirror actual real-world data (PCAPs). It also holds the CICFlowMeter network traffic investigation outcomes, which categorises flows based on the time brand, source and endpoint IP addresses, source and endpoint ports, protocols, and attack. Additionally, a definition of extracted characteristics is available. Using the HTTP, HTTPS, FTP, SSH, and email protocols, researchers constructed the abstract activities of 25 individuals for this dataset. Brute Force FTP, Brute Force SSH, DoS, Heartbleed, Web Attack, Infiltration, Botnet, and DDoS are some of the attacks used. They were executed on Tuesday, Wednesday, Thursday, and Friday in the morning and the afternoon. Sharafaldin et al. (2018). Database source: This data was generated by Sharafaldin et al. (2018). Specifically, the RF classifier, LR, and XGBoost have been examined among the ML systems in this study.
4.1 Experimental Goal
To the best of our knowledge, nobody has executed the following experiment,
  1. 1.
    Ensemble Stacking Classifier, where base classifiers are RF and XGBoost, the final LR classifier has never been implemented in SDN-based VANET as an IDS model. To gain optimum accuracy, FP/FN score and quick operation. And,
  2. 2.
    Bayesian Gaussian Hyperparameter Optimization to get the best parameter specs. And,
  3. 3.
    SelectKBest with chi² statistical test-based feature scoring contributed to this study’s best 10 feature selections.
  4. 4.
    The model’s performance was measured for the two datasets stated using Accuracy, Precision, Recall, F1, RMSE, and r2 Score, as well as the run-time of the IDS.
Finally, the above model will ensure the robustness and resilience of the proposed IDS model.
4.1.1 Virtual Machine Platform
To achieve the best possible hardware support for running the code, Microsoft Azure ML Studio Virtual Machine (VM) instance Standard E4s v3 (4 cores, 32 GB RAM, 64 GB HDD) and Processing unit CPU - Memory optimised were utilised for this experiment. The experiments for this research were carried out using Python 3.8.5, numpy Version 1.21.6, Pandas 1.1.5 and Scikit-Learn 0.23.1.
4.1.2 Supervised ML Approaches
Specifically, the RF classifier, LR, and XGBoost have been examined among the ML systems in this study.

### v5

@@ 3.1.3 Methods yet to reach optimum accuracy
According to findings of Mejri and Ben-Othman (2014), the proposed solution is a detection system that employs a statistical method, specifically linear regression, along with a watchdog mechanism to passively identify greedy behaviour at the MAC layer. A watchdog monitors the degree to which active nodes’ access periods correlate as part of the suggested approach. The network is under assault if the correlation coefficient is close to 1 but not precisely 1, and the linear regression straight-line slope is also not close to 1. This can happen when the correlation coefficient is near one but not quite one or when the slope is very close to 1 but not quite 1. The developed program does this by monitoring three metrics: the interval between transmissions, the total time spent transferring data, and the number of connect attempts made by a node. Subsequently, the performance of the solutions is analyzed in real-time using a trace file generated by SUMO. The SUMO scenario is constructed based on a city map that includes labels and traffic signals. To validate the reliability of the linear regression method, researchers initially simulate the network with nodes exhibiting correct behaviour. After that, four greedy automobiles are produced by injecting greedy nodes one at a time. To detect greedy behaviour, this approach takes 1.3 seconds for 1 node, 1.9 seconds for 2 nodes, 3.1 seconds for 3 nodes, and 7.9 seconds for 4 nodes. Ali Alheeti et al. (2015) The solution proposed in the paper is an ANN-based ill-use and anomaly intrusion detection system that can identify DoS assaults. The authors designed the solution by first developing a mobility scenario within SUMO. For ns-2 to read the files SUMO created, a conversion was performed using Mobility Vehicle (MOVE). In the end, ns-2 was utilised to imitate the vehicles’ communication ability. When considering the mobility and traffic situation, the authors used the urban mobility of Manhattan as their inspiration. The simulation takes two minutes and fifty seconds, and it involves cars and six Road roadside units (RSUs). A CBR (Constant Bit Rate) traffic that conveys User Datagram Protocol (UDP) packets is being performed by the vehicles. Only one of the automobiles intentionally tries to cause harm. It will discard the packets rather than send them to the destination. The solution built was accurate 85.02% of the time when classifying normal behaviours and 98.45% when classifying deviant behaviours.
@@
The authors of Gruebler et al. (2015) advocate using a Proportional Overlapping Scores (POS) approach to reduce the number of features taken from the trace file. These are fed into an artificial neural network (ANN) to teach it to classify data. After that, the ns-2 trace files that were created were employed in the POS algorithms so that the appropriate characteristics could be extracted. Next, Fuzzification is applied to the dataset’s data to prevent classification difficulties. Finally, the dataset is classified with the help of a Feed-Forward Neural Network (FFNN) using the IDS created for this solution. Sixty thousand records from the dataset were used, with half designated for training, another quarter for testing, and the remaining quarter for validation (25%). This dataset includes both benign and harmful activity in equal measure. The malicious behaviour vehicle was designed as a Black Hole attack, in which the malicious vehicle would not forward any received packets but dump them instead. Ultimately, the intrusion detection system was evaluated using anomaly and misuse detection. Regarding identifying misuse, it acquired a classification of 99.89 % for normal conduct and 99.80% for deviant behaviour. Regarding detecting anomalies, the findings for typical and aberrant behaviours came in at 99.87 and 99.72%, respectively.
@@
In their paper, Sedjelmaci and Senouci (2015) introduce a cluster-based intrusion detection system (IDS) to defend a network against selective forwarding, black holes, wormholes, packet repetition, resource depletion, and Sybil attacks. The technique that has been presented uses several different detection managers that operate on three different levels: cluster associate, cluster head, and RSU. The cars’ speeds determine how they should be organised into clusters. Cluster connection and the assurance of data safety are the criteria utilised in selecting the cluster heads. When social behaviour is introduced, the connectivity inside a cluster is improved. The IDS architecture comprises two primary detection schemes and a decision scheme. Local Intrusion Detection System and Global Intrusion Detection System are the detection systems run at cluster members and heads. The name of the decision-making scheme is Global Decision System, which is hosted on the RSUs. Because of this, the system can monitor several organisations and identify threats on multiple levels. In addition, each level can perform a unique set of algorithms and detecting approaches, which evaluate a unique set of characteristics. The Global Decision System will provide the collective standing of each car, which the cluster head will then send. The Worldwide Decision Scheme will then calculate the trust level of every vehicle. To evaluate the proposed system, ns-3 was used as the network simulator, while SUMO was used to mimic the vehicle’s movement. The IDS was evaluated concerning the detection rate, the false positive rate, and the detection time. The various assaults were put through their paces using a range of vehicle counts during testing. The percentage of hostile cars always equals 45% of the total. The provided findings range from 92% to 100%, although this depends on the total number of vehicles and the type of attack conducted.
@@
In their research, Ali Alheeti et al. (2016) propose an FFNN and SVM-based intrusion detection system is suggested in this paper. A methodical reaction is also meant to protect automobiles if harmful conduct is identified. A dataset composed of SUMO and ns-2 is utilised to train the IDS, and the features extracted from the trace file are used as training data. The POS algorithm has reduced these characteristics from 21 to 15. The collection consists of 30,000 entries that differentiate between appropriate and inappropriate behaviours. Fuzzification is performed on this dataset before it is used for classification. The writers separate the dataset into three categories: validation, test, and training. Generating grey-hole attacks involves picking malicious vehicles that drop packets at arbitrary intervals throughout the attack. The Ad Hoc On-Demand Distance Vector (AODV) protocol needs to be modified so that rushing assault can be carried out successfully. In SVM, the accuracy values achieved through simulation were 99.93% for normal behaviour and 99.64% for aberrant conduct. When FFNN was used, the findings showed an accuracy of 99.82% for normal behaviour and 98.86% accuracy for deviant conduct, respectively.
@@ 25 Features:
Proposed Framework
RF, CatBoost,
LR
VeReMi
Various attacks
(Constant, Random,
Eventual Stop)
Stacking
Ensemble with
SHAP Analysis
Accuracy: 99.68%,
Precision: 99.6%,
Recall: 99.31%,
F1-Score: 99.45%
@@
In their work, Alheeti et al. (2017) using Linear and Quadratic Discriminant Analysis, the authors present an intelligent IDS to guard against assaults, namely DoS and Black Hole attacks. To begin, evil actions are modelled and then emulated. Next, the authors make several adjustments to the AODV protocol to produce the DoS attack. In this scenario, denial-of-service attacks are carried out by dropping packets. After that, multiple mobilities and traffic setups were developed. SUMO and MOVE produced a realistic setting of both abnormal and typical behaviour. In addition to that, ns-2 was utilised to imitate various forms of communication. The dataset is derived from the ns-2 trace file, and all 21 characteristics are preserved throughout the process. The data is given a fuzzy appearance before being put through the test and the training. The findings of the Linear Discriminant Analysis indicate a detection rate of 86.44%, whereas the Quadratic Discriminant Analysis results indicate a detection rate of 85.67%.
@@
In the study conducted by Sharma and Kaul (2018), the authors suggest a novel technique for selecting a steady Cluster Head, and it uses Hybrid Fuzzy Multi-standards. After that, an ML-based intrusion detection system that uses SVM is utilised to identify malicious activities. The utilisation of a Dolphin Swarm Algorithm results in an improvement to the SVM-based IDS detection capabilities. This program uses dolphin swarms’ hunting and preying behaviour to locate and isolate misbehaving nodes in a network. The future method is evaluated through the use of simulation. Matlab and NetSim simulate the network, while SUMO models the traffic. During the simulation, a range of node densities from 50 to 300 are examined, with a maximum of 45% of the cars acting as attackers. The suggested intrusion detection system has a detection rate greater than 98%, selective forwarding, and wormhole attack; however, this percentage can vary based on the number of vehicles.
@@
As outlined by Kumar et al. (2015), a cloud intrusion detection system founded on clustering, and Learning Automate (LA) was presented. This method can potentially improve detection activity. The LA determines the leadership of clusters based on factors such as the total comparative velocity and the connectedness grade. The next stage, safeguarding the data, will begin after this procedure. In this process stage, an automaton will validate messages using the HMAC algorithm. When utilising the method that has been suggested, around 93% of harmful actions are recognised, and when mobility is added to leadership formation, the percentage of false positives is reduced. In addition to that, this technique is flexible enough to adapt to any alterations made to the network’s nodes.
@@
The work of Shu et al. (2021) presented an IDS for VANETs by implementing a Software-Defined Networking (SDN)-Based controller on each base station to differentiate between regular network traffic and attack network flows. Using the total network flow data, they trained several SDN controllers for the full VANET using GAN. This IDS approach permits distributed SDN controllers to detect their sub-network flows directly, and it can reduce high processing overheads. The author tested their algorithm using criteria along with precision, Accuracy, F1-Score, recall, and area under the curve (AUC), where maximum accuracy is 98.4%.
@@
As described by Zeng et al. (2019b), a deep learning-based intrusion detection method incorporating CNN and LSTM. The proposed model’s efficiency was compared to previously developed ML methods using a public dataset and an NS3 simulated VANET dataset with 28236 entries. The recommended model earned an F1 score ranging from 91% to 99%.
@@
In their analysis, Gad et al. (2021) discussed ML-based IDS techniques in VANET that are impressive. Researchers used SMOTE, chi2, and SMOTE+chi2 to eliminate the database imbalance problem in the feature engineering phase, and the XGBoost method’s overall score is better than others. In the Training-Testing phase, researchers used Logistic regression (LR), naive Bayes (NB), decision tree (DT), support vector machine (SVM), k-nearest neighbour (kNN), Random Forest (RF), AdaBoost, and XGBoost. This experiment could be better by using the Ensemble Sacking classification to increase overall performance, which this paper wants to address.
In the research conducted by Türkoğlu et al. (2022), the Minimum Redundancy Maximum Relevance (MRMR) feature selection algorithm to get the most correlated features for attack detection. First, scholars split the dataset in 3 ways: Training, Validation, and Test Data. After that, researchers used Bayesian Hyperparameter Optimization with training and validation data. However, 3-way splitting (60%, 20% and 20%) was not mandatory for tuning. Instead, the same training data could be used for tuning purposes. That way, 70% or 80% of data could be used for training the database, and 30% or 20% of data will be used for testing purposes to increase the overall accuracy.
As explored by Aneja et al. (2018), their research uses ANN (feed-forward network) with backpropagation technique. This backpropagation operates by scheming the gradient at every layer, known as the local gradient. Then, the weights and bias values are updated according to Levenberg-Marquardt optimisation. Moreover, a Genetic algorithm is used for feature selection.
In the study conducted by Alsarhan et al. (2021), the researchers introduce an optimised Support Vector Machine (SVM)-based Intrusion Detection System (IDS) for Vehicular Ad Hoc Networks (VANETs), employing Genetic Algorithm (GA), Particle Swarm Optimisation (PSO) and Ant Colony Optimization (ACO) to enhance accuracy. GA outperformed other algorithms, demonstrating improved detection rates and minimised false positives. A comprehensive evaluation validates the approach on the NSL-KDD dataset, providing significant insights into VANET security. Table 2 provides a summary of the key works and findings in relation to ML-based IDS models in VANET and SDVN.
@@
3.2 Advancing SD-VANETs with Explainable AI
Recent advancements in IDS tailored for SD-VANETs have leveraged a range of machine learning (ML) models. However, few of these systems have incorporated XAI to clarify their predictions. Prominent approaches, such as CNNs and RNNs, have been employed for anomaly detection (Zeng et al. (2019a)). However, these models frequently lack the transparency required for use in safety-critical environments, which can significantly diminish their effectiveness.
Emerging research has underscored the potential of XAI methodologies, particularly federated learning. This approach facilitates collaborative model training while ensuring data privacy and interoperability (Huang et al. (2023)). By utilizing decentralized data, federated learning not only boosts model accuracy but also sheds light on individual model contributions and decision-making processes, making it particularly relevant for the cybersecurity challenges faced by SD-VANETs.
In response to these challenges, our study seeks to enhance model interpretability by integrating SHAP analysis. This method clarifies feature importance and improves the transparency of model predictions. Additionally, we tackle persistent issues in IDS, such as elevated false positive rates and the identification of unknown attacks, by employing a stacked ensemble learning strategy. This approach has proven effective in enhancing overall performance through increased feature diversity and optimized decision boundaries, as evidenced in recent research by (Wang et al. (2024) and Kim et al. (2022)).
By merging SHAP analysis with ensemble techniques and leveraging the advantages of federated learning, our framework aims to deliver robust security solutions for SD-VANETs while promoting trust and accountability in automated decision-making processes.
@@ 4 Technical Implementation
The VeReMi dataset was created to test the effectiveness of VANET misbehaviour detection systems (vehicular networks). This database contains onboard unit message logs created from a simulation environment and tagged ground truth. The collection contains malicious messages that are meant to cause erroneous application behaviour, which misbehaviour detection techniques are designed to stop. The initial dataset includes five types of position falsification attacks. This data was generated by van der Heijden et al. (2018), and It has been collected from Sharma (2021).
To achieve the best possible hardware support for running the code, Microsoft Azure ML Studio Virtual Machine (VM) instance Standard E4s v3 (4 cores, 32 GB RAM, 64 GB HDD) and Processing unit CPU - Memory optimised were utilised for this experiment.
4.1 Supervised ML Approaches
Specifically, the RF classifier, LR, and CatBoost have been examined among the ML systems in this study.